OpenStudy (anonymous):

In the following reaction, when a surplus of hydrazine (N2H4) reacts with 20 liters of hydrogen peroxide (H2O2), how many liters of nitrogen gas will be released, given that both gases are at STP? N2H4 + 2H2O2 dN2 + 4H2O A) 10 liters B) 0.1 liters C) 40 liters D) 20 liters

OpenStudy (anonymous):

@thomaster or @chmvijay. help please stoichiometry =/

OpenStudy (frostbite):

Try write up the relative quantities you need for the reaction to run 1 mol times.

OpenStudy (anonymous):

thank you, i got the awnswer wrong but I think I know what i did wrong @Frostbite

OpenStudy (frostbite):

In that case explain to me what :)

OpenStudy (frostbite):

If you like?

OpenStudy (frostbite):

else in the future this is a relation that can help you if ever in doubt: for a reaction a A + b B -> c C d D It follows: \[\frac{ n(A) }{ a }=\frac{ n(B) }{ b }=\frac{ n(C) }{ c }=\frac{ n(D) }{ d }\] doing so we get: \[\frac{ n(N _{2}) }{ 1 }=\frac{ n(H _{2}O _{2} )}{ 2 }\] \[V(N _{2})=\frac{ V(H _{2}O _{2}) }{ 2 }\]